THEORETICAL PREDICTION OF THE ROTATIONAL CONSTANTS FOR PROTONATED METHANOL (CH_{3}OH_{2}^{+}): A MISSING PLAYER IN HOT CORE CHEMISTRY

\underline{D. E. WOON}, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana IL 61801.

Protonated molecules play important roles in the ion-molecule chemistry of hot cores. Protonated methanol, CH_{3}OH_{2}^{+}, is thought to contribute significantly to the gas phase formation of other known organic species such as methyl formate and dimethyl ether. While the protonated forms of many of the most common astromolecules have previously been detected, CH_{3}OH_{2}^{+} remains unidentified to date due to the absence of an experimental rotational spectrum. In order to provide theoretical predictions to assist anticipated new laboratory work, various levels of quantum chemical methodology were employed to predict the structure of CH_{3}OH_{2}^{+} and to convert its equilibrium rotational constants to A_{0}, B_{0}, and C_{0}. Potentials for the inversion and rotation of OH_{2}^{+} with respect to the CH_{3} group were also determined. This work built upon a benchmark study of methylamine discussed elsewhere (INSERT CODE).